### v2

# Closed-Form Approximation of the Total Variation Proximal Operator

Abstract
Total variation (TV) is a widely used function for regularizing imaging inverse problems that is particularly appropriate for images whose underlying structure is piecewise constant. TV regularized optimization problems are typically solved using proximal methods, but the way in which they are applied is constrained by the absence of a closed-form expression for the proximal operator of the TV function. A closed-form approximation of the TV proximal operator has previously been proposed, but its accuracy was not theoretically explored in detail. We address this gap by making several new theoretical contributions, proving that the approximation leads to a proximal operator of some convex function, it is equivalent to a gradient descent step on a smoothed version of TV, and that its error can be fully characterized and controlled with its scaling parameter. We experimentally validate our theoretical results on image denoising and sparse-view computed tomography (CT) image reconstruction.
This research was supported by the Laboratory Directed Research and Development program of Los Alamos National Laboratory under project numbers 20200061DR and 20230771DI.
1 Introduction
Inverse problems are typically posed as the estimation of a signal ${\bm{x}}\in\mathbb{R}^{n}$ from measurements
${\bm{y}}={\bm{A}}{\bm{x}}+{\bm{e}}\;,$ (1)
where ${\bm{A}}\in\mathbb{R}^{m\times n}$ denotes the forward operator and ${\bm{e}}\in\mathbb{R}^{m}$ represents additive noise. The classical approach to solving such problems is to pose them as an optimization problem
${\bm{x}}^{\ast}\in\mathop{\mathsf{arg\,min}}_{{\bm{x}}\in\mathbb{R}^{n}}f({\bm{x}})\quad\text{with}\quad f({\bm{x}})=g({\bm{x}})+\lambda h({\bm{x}})\;,$ (2)
where $g$ and $h$ represent the data-fidelity and regularization terms, respectively, and $\lambda\geq 0$ is the regularization parameter. When the noise in (1) corresponds to the additive white Gaussian noise (AWGN), it is common to use the squared $\ell_{2}$ data fidelity term $g({\bm{x}})=\frac{1}{2}\|{\bm{y}}-{\bm{A}}{\bm{x}}\|_{2}^{2}$. One of the most widely used choices for $h$ is total variation (TV) [1], which promotes solutions with a sparse image gradient. While it is no longer competitive for natural imagery, it remains effective for other types of imagery, such as those that occur in scientific or industrial imaging, where the implicit piecewise-constant image model is appropriate.
In this work, we focus on the two most common forms of TV, anisotropic $h^{\text{a}}$ and isotropic $h^{\text{i}}$, which are defined as
$h^{\text{a}}({\bm{x}})$ $:=\sum_{i=1}^{n}\sum_{j=1}^{d}\bigr{|}[{\bm{D}}_{j}{\bm{x}}]_{i}\bigr{|}$ (3)
$h^{\text{i}}({\bm{x}})$ $:=\sum_{i=1}^{n}\sqrt{\sum_{j=1}^{d}\bigr{(}[{\bm{D}}_{j}{\bm{x}}]_{i}\bigr{)}^{2}},$ (4)
where ${\bm{D}}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{nd}$ is the discrete gradient operator, $n$ denotes the signal size (e.g. $n=1024$ in a $32\times 32$ image), and $d$ represents the signal dimension (e.g. $d=2$ for images).
Gradient descent and related methods are not appropriate for problem (2) when $h$ is the TV function since it is non-smooth, but proximal methods [2] are very effective. They are based on the proximal operator of a function $h$, defined as
${\mathsf{prox}}_{\tau h}({\bm{z}}):=\mathop{\mathsf{arg\,min}}_{{\bm{x}}\in\mathbb{R}^{n}}\Bigg{\{}\frac{1}{2}\|{\bm{x}}-{\bm{z}}\|_{2}^{2}+\tau h({\bm{x}})\Bigg{\}},$ (5)
where $\tau>0$ is the proximal scaling parameter. This family of optimization algorithms includes the accelerated proximal gradient method (APGM) [3] and the alternating direction method of multipliers (ADMM) [4].
Unfortunately, there is no closed-form solution to the proximal operator (5) of the TV function, requiring it to be computed using an iterative approach [5] within algorithms such as APGM. These sub-iterations can be avoided via an appropriate ADMM variable splitting strategy [6], but in many cases this just pushes the computational complexity to a different component of the algorithm.
A closed-form approximation to the TV proximal operators was introduced in [7, 8, 9, 10], based on the concept of proximal average [11]. While the effectiveness of this approximation was analyzed within the context of APGM, its more general properties as an approximation to the TV proximal operator were not explored in detail, and its application in other proximal algorithms was not considered. In this paper, we address this gap by presenting new theoretical results demonstrating that the operator is indeed a good approximation to the proximal operators of anisotropic and isotropic TV. More specifically, we present two new contributions:
  1. (1)
    We provide new theoretical justifications for the approximation in [7, 8, 9, 10]. Specifically, we show that it is the proximal operator of some convex function, that it always decreases the TV function, and that its accuracy depends on the scaling parameter of the approximate proximal operator.
  2. (2)
    We provide new numerical results that validate our theoretical analysis by using the approximate TV proximal operator within APGM and ADMM algorithms for limited angle computed tomography reconstruction. These numerical results confirm the practical applicability of the approximate proximal operator and are consistent with the theory connecting the approximation accuracy with the scaling parameter.
2 Related Work
The TV regularizer promotes sparse image gradients, producing reconstructed images that are approximately piecewise constant. TV regularization has been demonstrated to be effective in various inverse problems, including image deconvolution, diffraction ultrasound tomography, compressed sensing, and optical tomography [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. APGM [3] is often employed to solve image reconstruction problems with the TV regularizer [5]. However, APGM requires additional sub-iterations to compute the proximal operator of the TV regularizer, which can slow down the overall convergence. Other approaches such as Primal-Dual Hybrid Gradient Algorithm (PDHG) [30] and ADMM [13] can be used to avoid these sub-iterations; but they have their own drawbacks such as increased memory consumption to represent the image gradients.
A fast non-iterative algorithm for solving the 1D TV proximal operator has been proposed [31], but it does not generalize for signals with more than one dimension, such as images. To generalize to an arbitrary number of dimensions, a closed-form approximation of the proximal operator for TV was proposed, eliminating the need for sub-iterations [7, 8, 10, 9]. The convergence of APGM with this approximation for anisotropic TV was investigated in [9] and for isotropic TV was investigated in [10]. These approximations share a close relationship with image reconstruction techniques using wavelet domain regularization [32]. They can also be viewed as the application of the proximal average [11] approximation for the proximal operator of a sum of multiple nonsmooth functions. Cycle spinning is another closely related concept originally introduced for denoising [33], later refined [34], and applied in various imaging inverse problems, including image restoration, MRI reconstruction, 3D CT reconstruction, and computed tomography [35, 36, 37, 38, 39, 40, 41, 42, 43].
The convergence of proximal algorithms has been extensively investigated [2]. Our work is more related to the family of inexact proximal algorithms that rely on approximations of proximal operators or gradient calculations [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. In particular, our theoretical analysis shows that the approximation in [7, 8, 10, 9] leads to an inexact TV proximal operator whose accuracy depends on its scaling parameter, thus enabling its use within existing proximal methods.
Algorithm 1 APGM with ${\mathcal{S}}_{\tau}({\bm{z}})$
0:$g$, ${\bm{x}}_{0},{\bm{z}}_{0},{\bm{s}}_{0}\in\mathbb{R}^{n}$, $\gamma>0$, $\lambda>0$, and $\{q_{k}\}_{k\in\mathbb{N}}$
1:for $k=1,2,$... do
2:  ${\bm{z}}^{k}={\bm{s}}^{k-1}-\gamma\nabla g({\bm{s}}^{k-1})$
3:  ${\bm{x}}^{k}={\mathcal{S}}_{\gamma\lambda}({\bm{z}}^{k})$
4:  ${\bm{s}}^{k}={\bm{x}}^{k}+((q_{k-1}-1)/q_{k})({\bm{x}}^{k}-{\bm{x}}^{k-1})$
5:end for
Algorithm 2 ADMM with ${\mathcal{S}}_{\tau}({\bm{z}})$
0:$g$, ${\bm{x}}_{0},{\bm{z}}_{0},{\bm{s}}_{0}\in\mathbb{R}^{n}$, $\gamma>0$, $\lambda>0$
1:for $k=1,2,$... do
2:  ${\bm{z}}^{k}={\mathsf{prox}}_{\gamma g}({\bm{x}}^{k-1}-{\bm{s}}^{k-1})$
3:  ${\bm{x}}^{k}={\mathcal{S}}_{\gamma\lambda}({\bm{z}}^{k-1}+{\bm{s}}^{k-1})$
4:  ${\bm{s}}^{k}={\bm{s}}^{k-1}+{\bm{x}}^{k}-{\bm{z}}^{k}$
5:end for
3 Approximating Total Variation
3.1 Method
In this section, we present the approximation of the TV proximal operator proposed in [9, 10]. The motivation for this approximation is the reduction of the computational complexity of calculating the TV proximal operator from $O(knd)$, where $k$ is the number of sub-iterations, to $O(nd)$. Consider the linear mapping ${\bm{W}}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{2nd}$ defined as
${\bm{W}}:=\frac{1}{2\sqrt{d}}\begin{bmatrix}{\bm{M}}\vskip 3.0pt\\ {\bm{D}}\end{bmatrix},\qquad\begin{aligned} {\bm{M}}&:=\big{[}{\bm{M}}_{1}^{{\mathsf{T}}}~\cdots~{\bm{M}}_{d}^{{\mathsf{T}}}\big{]}^{\mathsf{T}},\\ {\bm{D}}&:=\big{[}{\bm{D}}_{1}^{{\mathsf{T}}}~\cdots~{\bm{D}}_{d}^{{\mathsf{T}}}\big{]}^{\mathsf{T}},\end{aligned}$
where ${\bm{M}}_{j}:\ \mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ for $j\in\{1,...,d\}$ represents an averaging operator along the $j^{\text{th}}$ dimension, and ${\bm{D}}_{j}:\ \mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ for $j\in\{1,...,d\}$ denotes a discrete gradient operator along the $j^{\text{th}}$ dimension. ${\bm{M}}_{j}$ and ${\bm{D}}_{j}$ are convolutions along the $j^{\text{th}}$ dimension of the signal with the kernels
${\bm{m}}=\begin{bmatrix}1\vskip 3.0pt\\ 1\end{bmatrix},\qquad{\bm{d}}=\begin{bmatrix}1\vskip 3.0pt\\ -1\end{bmatrix}$,
respectively. In this work, we assume periodic boundary conditions for both $\bm{M}$} and $\bm{D}$}. Note that this is not essential for the method or the results in this paper: aperiodic boundary conditions are perfectly acceptable. The mapping $\bm{W}$} can be interpreted as a union of scaled and shifted first-level discrete Haar wavelet and scaling functions across all $d$ dimensions [9]. It is important to note that this operator satisfies ${\bm{W}}^{{\mathsf{T}}}{\bm{W}}=\bm{I}$, but due to redundancy in the union of scaled orthogonal transforms, ${\bm{W}}{\bm{W}}^{{\mathsf{T}}}$ is over-determined and so ${\bm{W}}{\bm{W}}^{{\mathsf{T}}}\neq\bm{I}$ [55] (Section 3, first paragraph).
The approximate TV proximal operator is defined as
${\mathcal{S}}_{\tau}({\bm{z}})\,\coloneqq\,{\bm{W}}^{{\mathsf{T}}}{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}}),$ (6)
where ${\mathcal{T}}_{\lambda}$ is an operator with the parameter $\lambda\,\coloneqq\,\tau\hskip 1.49994pt2\sqrt{d}>0$, corresponding to the application of component-wise shrinkage functions on the scaled difference components of ${\bm{W}}{\bm{z}}$. For the anisotropic operator ${\mathcal{T}}^{\hskip 1.22911pta}_{\tau 2\sqrt{d}}$, the shrinkage function
${\mathsf{max}}(|u|-\lambda,0)\frac{u}{|u|},$ (7)
is applied to each difference component; that is, $u=[{\bm{W}}{\bm{z}}]^{\text{dif}}_{i}\in\mathbb{R}$ for $i\in\{1,2,...,nd\}$. For the isotropic operator ${\mathcal{T}}^{\hskip 1.22911pti}_{\tau 2\sqrt{d}}$, the shrinkage function
${\mathsf{max}}(\|{\bm{u}}\|_{2}-\lambda,0)\frac{{\bm{u}}}{\|{\bm{u}}\|_{2}},$ (8)
is applied to each finite difference vector corresponding to each pixel in $\bm{z}$}; that is, ${\bm{u}}=[{\bm{W}}{\bm{z}}]^{\text{dif}}_{G_{i}}\in\mathbb{R}^{d}$ for $i\in\{1,2,\cdots,n\}$. Each $G_{i}$ contains the $d$ indices in ${\bm{W}}{\bm{z}}]^{\text{dif}}$ corresponding to the components of the finite difference vector at pixel $i$.
Note that the shrinkage parameter $\tau$ must be scaled by $2\sqrt{d}$ due to the scaling of $\bm{W}$}. For simplicity, we omit the superscripts a and i when we do not need to distinguish between the anisotropic and isotropic thresholding functions, and the corresponding TV functions. The following results apply to both versions.
3.2 Theoretical Analysis
We now present the theoretical analysis ${\mathcal{S}}_{\tau}({\bm{z}})$ as an approximation of the TV proximal operator. The detailed proofs are provided in the Appendix 6.
Proposition 1.
For $\tau>0$, there exists a proper, closed, and convex function $\widehat{h}$ such that ${\mathcal{S}}_{\tau}({\bm{z}})={\mathsf{prox}}_{\widehat{h}}({\bm{z}})$ for all ${\bm{z}}\in\mathbb{R}^{n}$.
Figure 1: Effect of $\tau$ on image denoising performance when using ${\mathcal{S}}_{\tau}({\bm{z}})$ in APGM compared to the exact TV reconstruction with regularization $\lambda=0.5$. As written in Algorithm 1, $\tau=\gamma\lambda$, where $\gamma$ is the step-size. The top-left shows the relative cost of the approximate TV reconstruction to the exact TV reconstruction. The PSNR and difference images are relative to the exact TV reconstruction. Following Proposition 3, a smaller $\tau$ results in a smaller error.
Proposition 1 establishes the existence of a function $\widehat{h}$ for which ${\mathcal{S}}_{\tau}({\bm{z}})$ is the proximal operator. Although the explicit form of $\widehat{h}$ is unknown, its existence implies that ${\mathcal{S}}_{\tau}({\bm{z}})$ can be incorporated into any proximal optimization method with the expectation of algorithm convergence.
We will next demonstrate that ${\mathcal{S}}_{\tau}({\bm{z}})$ provides a close approximation of the proximal operator of TV. First, we show that it is related to the following smoothed versions of $h$. In the anisotropic case, define
$\widetilde{h}^{a}({\bm{z}})=\sum_{i=1}^{nd}\phi_{\tau 4d}([{\bm{D}}{\bm{z}}]_{i}),$ (9)
where $\phi_{\theta}:\mathbb{R}\rightarrow\mathbb{R}$ is defined as
$\phi_{\theta}(y)=\begin{cases}\frac{y^{2}}{2\theta}&\text{for }|y|\leq\theta\\ |y|-\frac{\theta}{2}&\text{for }|y|>\theta\end{cases}.$ (10)
For the isotropic case, define
$\widetilde{h}^{i}({\bm{z}})=\sum_{i=1}^{n}\psi_{\tau 4d}([{\bm{D}}{\bm{z}}]_{G_{i}}),$ (11)
where for ${\bm{y}}\in\mathbb{R}^{d}$ we define
$\psi_{\theta}({\bm{y}})=\begin{cases}\frac{\|{\bm{y}}\|^{2}}{2\theta}&\text{for }\|{\bm{y}}\|\leq\theta\\ \|{\bm{y}}\|-\frac{\theta}{2}&\text{for }\|{\bm{y}}\|>\theta\end{cases}.$ (12)
Note that each $G_{i}$ is the collection of $d$ indices of the vector ${\bm{D}}{\bm{z}}$ corresponding to the finite difference vector at position $i$ of $\bm{z}$}. Related approximations have previously appeared in the TV literature [30, 56, 57, 58, 59, 60], based on the Huber function [61] or Moreau-Yoshida regularization [62, 63].
Proposition 2.
Let $\widetilde{h}^{a}({\bm{z}})$ and $\widetilde{h}^{i}({\bm{z}})$ denote the Huber variants of anisotropic TV and isotropic TV given in Eq. (9) and Eq. (11), respectively. Then,
${\mathcal{S}}^{a}_{\tau}({\bm{z}})={\bm{z}}-\tau\nabla\widetilde{h}^{a}({\bm{z}})\quad\text{and}\quad{\mathcal{S}}^{i}_{\tau}({\bm{z}})={\bm{z}}-\tau\nabla\widetilde{h}^{i}({\bm{z}}).$
Proposition 2 states that the operator ${\mathcal{S}}_{\tau}({\bm{z}})$ is equivalent to a gradient step with step-size $\tau$ on a smoothed version of TV. A recent work [64] demonstrates that running a single iteration of the dual forward backward algorithm to solve the TV proximal within a FPG algorithm results in an algorithm that converges to a solution of the TV minimized problem. Proposition 2 shows that ${\mathcal{S}}_{\tau}({\bm{z}})$ is related; however, instead of running one iteration of the dual-FB algorithm, ${\mathcal{S}}_{\tau}({\bm{z}})$ corresponds to one iteration of gradient descent on a smoothed version of TV. To decrease Eq. (9) or Eq. (11), the step-size must be at most $1/L$, where $L$ is the Lipschitz constant of the function’s gradient [65] (Section 9.3). For both Eq. (9) and Eq. (11), the Lipschitz constant is $1/{\tau}$, as verified in the Appendix, satisfying the condition.
To further describe the operator as an approximation of the true proximal operator for h{h}, we introduce the concept of the subdifferential [62, 66].
Definition 1.
The subdifferential of the convex function $f:\mathbb{R}^{n}\rightarrow(-\infty,+\infty]$ at ${\bm{y}}\in\text{dom}(f)$ is defined to be
$\partial f({\bm{y}}):=\big{\{}{\bm{g}}\in\mathbb{R}^{n}:f({\bm{x}})\geq f({\bm{y}})+{\bm{g}}^{{\mathsf{T}}}({\bm{x}}-{\bm{y}}),\hskip 1.49994pt\forall{\bm{x}}\in\mathbb{R}^{n}\big{\}}.$
A relaxation of this definition is called the $\epsilon$-subdifferential [67] and defined as
Definition 2.
The $\epsilon$-subdifferential of the convex function $f:\mathbb{R}^{n}\rightarrow(-\infty,+\infty]$ at ${\bm{y}}\in\text{dom}(f)$ is defined to be
$\partial_{\epsilon}$ $f({\bm{y}})\,\coloneqq\,\big{\{}{\bm{g}}\in\mathbb{R}^{n}:f({\bm{x}})\geq f({\bm{y}})+{\bm{g}}^{{\mathsf{T}}}({\bm{x}}-{\bm{y}})-\epsilon,\forall{\bm{x}}\in\mathbb{R}^{n}\big{\}}.$
By comparing Definition 2 to Definition 1, it becomes evident that the traditional subdifferential is a subset of the $\epsilon$-subdifferential. The $\epsilon$-subdifferential represents the subdifferentials relaxed by $\epsilon$.
Substantial work has been done to define and analyze various notions of approximate proximal operators [44, 51, 52, 49]. In Proposition 3, we prove that the operator ${\mathcal{S}}_{\tau}({\bm{z}})$ can be characterized using several of the common definitions for an approximate proximal operator [49].
Proposition 3.
For all ${\bm{z}}\in\mathbb{R}^{n}$, the operator ${\mathcal{S}}_{\tau}({\bm{z}})$ satisfies
  1. (a)
    ${\mathcal{S}}_{\tau}({\bm{z}})={\mathsf{prox}}_{\tau{h}}({\bm{z}}+{\bm{\delta}})$, where $\|{\bm{\delta}}\|_{2}\leq\tau\epsilon_{1}$,
  2. (b)
    ${\bm{z}}-{\mathcal{S}}_{\tau}({\bm{z}})\in\tau\partial_{\tau\epsilon_{2}}{h}({\mathcal{S}}_{\tau}({\bm{z}}))$,
where $\epsilon_{1}$ and $\epsilon_{2}$ are constants.
Proposition 3 relies on notions of approximate proximal operators, commonly used in the literature. Specifically, Proposition3 (b) aligns with a “type 3 approximation” in the proximal operator with $\tau\epsilon_{1}$-precision, while Proposition 3 (c) corresponds to a “type 2 approximation” with $\tau\sqrt{2\epsilon_{2}}$-precision [49].
The first part of Proposition 3 demonstrates that the operator functions as a perturbed proximal operator of TV. Specifically, the output of the operator is equivalent to applying the true proximal operator of TV to a perturbed input vector $\bm{z}$}, modified by a vector $\bm{\delta}$}. Notably, the norm of the perturbation $\|{\bm{\delta}}\|_{2}$ is bounded and can be controlled by adjusting $\tau$. From the optimality condition for the proximal operator of true TV regularizer ${\mathsf{prox}}_{\tau{h}}({\bm{z}})$, there is the following well known relationship
${\bm{z}}-{\mathsf{prox}}_{\tau{h}}({\bm{z}})\in\tau\partial\hskip 1.49994pt{h}({\mathsf{prox}}_{\tau{h}}({\bm{z}}))\;.$
The second part of Proposition 3 establishes an analogous connection for the closed-form approximate operator, characterized using the $\epsilon$-subdifferential from Definition 2, where $\epsilon$ is controlled by the proximal scaling parameter $\tau$.
Overall, Proposition 3 implies that the $\widehat{h}$ from Proposition 1 can be seen as an approximation of the true TV function h{h} with controllable accuracy. Notably, by reducing $\tau$, one can achieve an arbitrarily close approximation to the true TV proximal operator. Therefore, our theoretical result supports the use of ${\mathcal{S}}_{\tau}({\bm{z}})$ as an approximation of ${\mathsf{prox}}_{\tau{h}}({\bm{z}})$. As the accuracy of the operator is controlled by $\tau$, it might initially appear that it would only be effective for problems with a very small regularization parameter $\lambda$ for TV. However, this is not the case when ${\mathcal{S}}_{\tau}({\bm{z}})$ is used within iterative algorithms such as APGM and ADMM. As shown in Algorithms 1 and 2, the shrinkage parameter for ${\mathcal{S}}_{\tau}({\bm{z}})$ is set by $\tau=\gamma\lambda$. Thus, one can enhance accuracy by adjusting $\gamma$ rather than altering the TV regularization parameter $\lambda$ . For APGM, $\gamma$ represents the step size, while for ADMM, $\gamma$ serves as the penalty parameter in the augmented Lagrangian.
Figure 2: Effect of $\tau$ on the approximate reconstruction performance $\widehat{f}$ using the ${\mathcal{S}}_{\tau}({\bm{z}})$ relative to the exact TV reconstruction $f^{*}$. ${\mathcal{S}}_{\tau}({\bm{z}})$ is tested within the APGM and ADMM algorithms. For image denoising $\lambda=0.5$ and for CT reconstruction $\lambda=5$. In both cases, as written in Algorithms 1 and 2, $\tau=\gamma\lambda$, where $\gamma$ is the step-size in APGM and the penalty parameter in ADMM. As expected from Proposition 3, the smaller the $\tau$, the smaller the error.
Table 1: The relative cost, PSNR with respect to the exact TV solution and groundtruth, and acceleration achieved by using ${\mathcal{S}}_{\tau}({\bm{z}})$ (measured as the ratio of the number of iterations when using ${\mathcal{S}}_{\tau}({\bm{z}})$ to total number of sub-iterations when using the exact TV proximal) for various regularization ($\lambda$) and step-size/penalty ($\gamma$) parameters for image denoising and CT reconstruction. For the APGM and ADMM algorithms, $\tau=\gamma\lambda$, where $\gamma$ determines the accuracy of ${\mathcal{S}}_{\tau}({\bm{z}})$. The metrics have been averaged across the 10 images in the dataset. Note that $L$ is the Lipschitz constant of $\nabla g$.
APGM Image Denoising
$\lambda=0.25$ $\lambda=0.5$ $\lambda=1$
Step Size ($\gamma$) Rel. Err. PSNR (TV) PSNR (GT) Accel Rel. Err. PSNR (TV) PSNR (GT) Accel Rel. Err. PSNR (TV) PSNR (GT) Accel
1e-1 1.751e-02 31.97 16.44 $2.32\times$ 1.214e-01 23.14 17.57 $9.58\times$ 3.875e-01 19.65 17.38 $19.86\times$
1e-2 1.364e-03 48.02 16.72 $0.41\times$ 1.281e-02 34.52 18.57 $1.61\times$ 5.467e-02 25.82 15.68 $2.67\times$
1e-3 1.152e-04 65.90 16.72 $0.09\times$ 1.157e-03 49.16 18.75 $0.38\times$ 5.945e-03 36.31 15.96 $0.67\times$
APGM CT Reconstruction
$\lambda=2.5$ $\lambda=5$ $\lambda=10$
Step Size ($\gamma$) Rel. Err. PSNR (TV) PSNR (GT) Accel Rel. Err. PSNR (TV) PSNR (GT) Accel Rel. Err. PSNR (TV) PSNR (GT) Accel
1/(1LL) 9.473e-04 57.28 21.07 $49.50\times$ 2.069e-03 52.08 20.86 $49.21\times$ 4.588e-03 46.48 20.41 $49.71\times$
1/(2LL) 4.609e-04 59.46 21.06 $39.99\times$ 1.002e-03 56.87 20.88 $40.20\times$ 2.208e-03 51.84 20.48 $39.21\times$
1/(4LL) 2.264e-04 62.14 21.07 $33.51\times$ 4.874e-04 60.42 20.88 $32.99\times$ 1.068e-03 56.79 20.51 $31.09\times$
ADMM CT Reconstruction
$\lambda=2.5$ $\lambda=5$ $\lambda=10$
Pen. Par. ($\gamma$) Rel. Err. PSNR (TV) PSNR (GT) Accel Rel. Err. PSNR (TV) PSNR (GT) Accel Rel. Err. PSNR (TV) PSNR (GT) Accel
1e-2 3.545e-01 18.49 15.35 $260.90\times$ 4.341e-01 18.10 15.13 $157.57\times$ 4.680e-01 18.28 15.13 $89.96\times$
1e-3 3.767e-02 31.63 19.82 $62.75\times$ 8.291e-02 26.42 18.59 $69.34\times$ 1.788e-01 21.91 16.82 $69.53\times$
1e-4 4.850e-03 40.42 20.59 $11.43\times$ 9.284e-03 38.55 20.37 $10.24\times$ 1.989e-02 35.13 19.82 $9.90\times$
4 Numerical Validation
In this section, we numerically validate Proposition 3 by comparing the reconstruction performance of iterative proximal algorithms when ${\mathcal{S}}_{\tau}({\bm{z}})$ replaces the true proximal operator for isotropic total variation. For experimental validation of the efficiency and timing of these operators, we refer the reader to [8, 9, 10]. Specifically, we run APGM and ADMM on a simulated low-angle computed tomography (CT) image reconstruction problem. The APGM and ADMM algorithms using ${\mathcal{S}}_{\tau}({\bm{z}})$ instead of the true proximal operator are shown in Algorithms 1 and 2, respectively. The accuracy of ${\mathcal{S}}_{\tau}({\bm{z}})$ can be directly controlled by $\gamma$ since $\tau=\gamma\lambda$, where $\gamma$ is the step size and penalty parameter, for APGM and ADMM algorithms, respectively. To calculate the true proximal operator, we use the iterative Fast Projected Gradient (FPG) algorithm [5]. Additionally, we investigate the performance of the closed-form approximate operator relative to true proximal operator of TV for an image denoising task. The numerical evaluations are performed using 10 piecewise-smooth foam phantoms generated by the XDesign Python package [68] and the implementationsSee classes AnisotropicTVNorm and IsotropicTVNorm. of ${\mathcal{S}}_{\tau}({\bm{z}})$ in the SCICO Python package [69]. All algorithms are run until the stopping criteria
$\frac{\|{\bm{x}}^{t}-{\bm{x}}^{t-1}\|_{2}}{\|{\bm{x}}^{t-1}\|_{2}}\leq 5\times 10^{-6}$ (13)
is met.
4.1 Image Denoising
In the case of image denoising, the forward model is ${\bm{A}}={\bm{I}}$, which results in the data-fidelity term $g({\bm{x}})=\frac{1}{2}\|{\bm{y}}-{\bm{x}}\|_{2}^{2}$. We report the denoising results for the regularization parameters $0.25$, $0.5$, and $1$ in Table 1. Since FPG solves the denoising problem, it is directly used to compute the exact TV regularized denoised image. The APGM algorithm with ${\mathcal{S}}_{\tau}({\bm{z}})$ is used to compute the approximate TV regularized images.
4.2 Computed Tomography
We validate the application of ${\mathcal{S}}_{\tau}({\bm{z}})$ for the limited-angle computed tomography (CT) reconstruction. The data fidelity term in Eq. (2) is $g({\bm{x}})=\frac{1}{2}\|{\bm{A}}{\bm{x}}-{\bm{y}}\|_{2}^{2},$ where $\bm{A}$} represents the CT imaging forward operator using only $45$ angles, resulting in an ill-posed inverse problem. We show the experimental results for the regularization parameters $\lambda$ corresponding to $2.5$, 55, and $10$, using both APGM and ADMM. ADMM is implemented to solve the constrained optimization problem
$\mathop{\mathsf{arg\,min}}_{{\bm{x}},{\bm{z}}}g({\bm{x}})+\lambda h({\bm{z}})\quad\text{s.t.}\;\;{\bm{z}}={\bm{x}}$
in order to test the impact of ${\mathcal{S}}_{\tau}({\bm{z}})$. The exact TV reconstruction is obtained by running 50 sub-iterations of FPG to calculate the exact proximal within both algorithms.
4.3 Discussion
According to Proposition 1, ${\mathcal{S}}_{\tau}({\bm{z}})$ serves as the proximal operator for some convex function. Consequently, proximal-based reconstruction algorithms using this operator are guaranteed to converge. Activation of the termination criteria in 13 occurs for all the experiments, validating the convergence of the algorithms under ${\mathcal{S}}_{\tau}({\bm{z}})$.
Figure 3: Effect of $\tau$ on limited angle Computed Tomography (CT) reconstruction using ${\mathcal{S}}_{\tau}({\bm{z}})$ in APGM compared to the exact TV reconstruction with regularization $\lambda=5$. As written in Algorithm 1, $\tau=\gamma\lambda$, where $\gamma$ is the step-size. The top-left shows relative cost of the approximate reconstruction to the exact reconstruction. The PSNR and difference images are relative to the exact reconstruction. Following Proposition 3, a smaller $\tau$ results in a smaller error.
The statements in Proposition 2 and Proposition 3 imply that a smaller $\tau$ in ${\mathcal{S}}_{\tau}({\bm{z}})$ provides a better approximation with respect to the true TV reconstruction. This relationship is experimentally validated in Table 1 across all experiments: a smaller $\tau$ results in a smaller relative error in the approximate reconstruction loss $\widehat{f}$, compared to the true TV reconstruction loss $f^{*}$. Relative error, denoted as Rel. Err. in Table 1, is defined to be
$\frac{\widehat{f}-f^{*}}{f^{*}}.$
Similarly, the PSNR with respect to the true TV reconstruction increases as $\tau$ decreases, suggesting ${\mathcal{S}}_{\tau}({\bm{z}})$ provides a better approximation of the TV proximal operator. Table 1 also displays the trade-off between approximation accuracy and acceleration when using ${\mathcal{S}}_{\tau}({\bm{z}})$ compared to $50$ iterations of FPG for the exact proximal. “Accel” is computed as the number of iteration using ${\mathcal{S}}_{\tau}({\bm{z}})$ divided by total number of sub-iterations when using FPG for the exact proximal. Figure 2 illustrates the dependence of cost accuracy on parameter $\gamma$ for image denoising with $\lambda=0.5$ and CT image reconstruction with $\lambda=5$. Figures 1,  3, and 4 illustrate visual comparison for image denoising and CT reconstruction using true TV proximal and ${\mathcal{S}}_{\tau}({\bm{z}})$ operator with various $\gamma$. The visual results indicate that ${\mathcal{S}}_{\tau}({\bm{z}})$ promotes piecewise smooth images, as is expected from an approximate TV proximal operator. Both visual and numerical results from the experiments support the theoretical finding presented in Section 3.2, demonstrating that ${\mathcal{S}}_{\tau}({\bm{z}})$ can approximate the TV proximal and be effectively integrated into proximal-based iterative algorithms to achieve approximate TV-regularized reconstructions.
Figure 4: Effect of $\tau$ on limited angle Computed Tomography (CT) reconstruction using ${\mathcal{S}}_{\tau}({\bm{z}})$ in ADMM compared to exact reconstruction with regularization $\lambda=5$. As written in Algorithm 2, $\tau=\gamma\lambda$, where $\gamma$ is the penalty parameter. The top-left shows relative cost of the approximate reconstruction to the exact reconstruction. The PSNR and difference images are relative to exact reconstruction. Following Proposition 3, a smaller $\tau$ results in a smaller error.
5 Conclusion
We theoretically analyzed a closed-form operator that approximates the proximal operator for both anisotropic and isotropic total variation. While these operators have been previously explored in the context of optimization algorithms, this work focuses on the operator itself. Specifically, we demonstrated that the operator consistently reduces the TV function of the input signal and that its error, relative to the true TV proximal operator, can be characterized using well-established notions of approximate proximal operators. Through this analysis, we provide a foundation for employing ${\mathcal{S}}_{\tau}({\bm{z}})$ within any proximal-based reconstruction method to address TV-regularized imaging inverse problems. Finally, we supported our theoretical findings with several experiments.
Acknowledgment
The authors thank the anonymous reviewers for their careful reading and comments that helped improve the final version of this manuscript.
6 Appendix
In this section, we will prove Propositions  12, and  3. The linear map ${\bm{W}}:R^{n}\rightarrow\mathbb{R}^{2nd}$ is defined as
${\bm{W}}:=\frac{1}{2\sqrt{d}}\begin{bmatrix}{\bm{M}}\vskip 3.0pt\\ {\bm{D}}\end{bmatrix},\qquad\begin{aligned} {\bm{M}}&:=\big{[}{\bm{M}}_{1}^{{\mathsf{T}}}~\cdots~{\bm{M}}_{d}^{{\mathsf{T}}}\big{]}^{\mathsf{T}},\\ {\bm{D}}&:=\big{[}{\bm{D}}_{1}^{{\mathsf{T}}}~\cdots~{\bm{D}}_{d}^{{\mathsf{T}}}\big{]}^{\mathsf{T}},\end{aligned}$
where $\bm{M}$} is the averaging map and $\bm{D}$} is the finite difference map. Therefore, for any vector ${\bm{u}}\in\mathbb{R}^{2nd}$ half of the components correspond to the averaging map and the remaining components correspond to the finite difference map. We use the notation ${\bm{u}}^{\text{dif}}$ to correspond to the difference coefficients.
Using this notation, we define the function $\bar{h}:\mathbb{R}^{2nd}\rightarrow\mathbb{R}$, where $\bar{h}$ has an anisotropic and isotropic form, $\bar{h}^{\text{a}}$ and $\bar{h}^{\text{i}}$, respectively. Let $G_{1},...,G_{n}$ be a partition of $\{1,2,...,nd\}$ such that each $G_{i}$ is the set of indices corresponding to the finite difference of pixel $i$. For ${\bm{u}}\in\mathbb{R}^{2nd}$,
$\bar{h}^{\text{a}}({\bm{u}})$ $:=2\sqrt{d}\|{\bm{u}}^{\text{dif}}\|_{1,1}=2\sqrt{d}\sum_{i=1}^{n}\|[{\bm{u}}^{\text{dif}}]_{G_{i}}\|_{1}$ (14)
$\bar{h}^{\text{i}}({\bm{u}})$ $:=2\sqrt{d}\|{\bm{u}}^{\text{dif}}\|_{2,1}=2\sqrt{d}\sum_{i=1}^{n}\|[{\bm{u}}^{\text{dif}}]_{G_{i}}\|_{2}\;,$ (15)
where $\|{\bm{u}}^{\text{dif}}\|_{p,1}$ is a group norm, with $p$ either $1$ or $2$, to anisotropic and isotropic TV, respectively. The grouping is based on the $d$ elements that represent the discrete differences at a specific location for each of the $d$ dimensions, with the $p$-norm applied within each group. Each ${\bm{u}}^{\text{dif}}]_{G_{i}}\in\mathbb{R}^{d}$ is the vector at the indices given by the set $G_{i}$. We use
$\bar{h}({\bm{u}}):=2\sqrt{d}\|{\bm{u}}^{\text{dif}}\|_{p,1}$ (16)
to refer to both the anisotropic and isotropic versions. Thus the results in Proposition 1 and Proposition 3 hold for both anisotropic and isotropic versions. We can also rewrite the TV definitions in Eq. (3) and Eq. (4) in the general form
${h}({\bm{z}})=\|{\bm{D}}{\bm{z}}\|_{p,1}\;.$ (17)
The term $2\sqrt{d}$ in front of $\bar{h}$ is necessary so that the relationship
$\bar{h}({\bm{W}}{\bm{z}})$ $={h}({\bm{z}})$ (18)
holds. Note that Eq. (18) implies that for ${\bm{z}}\in\mathbb{R}^{n}$, $\bar{h}$ evaluated at ${\bm{W}}{\bm{z}}$ is equivalent to h{h} evaluated at $\bm{z}$}; however, $\bar{h}$ is not the TV function itself since it is defined on the entire $\mathbb{R}^{2nd}$, and not $\mathbb{R}^{n}$.
Additionally, $\bar{h}$ is defined as the scaled version of two well-known norms that are proper, closed, and convex. By using the notation $\Gamma^{0}$ to represent the set of proper, closed, and convex functions, we have $\bar{h}\in\Gamma^{0}(\mathbb{R}^{2nd})$. Finally, define the linear subspace ${\mathcal{U}}:=\{{\bm{u}}|\ {\bm{u}}={\bm{W}}{\bm{W}}^{{\mathsf{T}}}{\bm{u}}\}\subset\mathbb{R}^{2nd}$.
Proof of Proposition 1
Proposition 1.
For $\tau>0$, there exists $\widehat{h}\in\Gamma^{0}(\mathbb{R}^{n})$ such that ${\mathcal{S}}_{\tau}({\bm{z}})={\mathsf{prox}}_{\widehat{h}}({\bm{z}})$ for all ${\bm{z}}\in\mathbb{R}^{n}$.
Proof.
To establish that ${\mathcal{S}}_{\tau}({\bm{z}})$ is a proximal operator of some $\widehat{h}\in\Gamma^{0}(\mathbb{R}^{n})$, it is sufficient to show that
  1. 1.
    There exists a closed and convex $\psi$ such that ${\mathcal{S}}_{\tau}({\bm{z}})\in\partial\psi({\bm{z}}),\ \forall{\bm{z}}\in\mathbb{R}^{n}$.
  2. 2.
    The operator ${\mathcal{S}}_{\tau}({\bm{z}})$ is nonexpansive.
This equivalence was given in Corollary 10.c in [62]. As defined in Eq. (16), $\bar{h}$ is the scaled norm of the difference components of its input vector in $\mathbb{R}^{2nd}$. Therefore, the proximal operator of $\tau\bar{h}({\bm{u}})$ can be written using the thresholding function defined in Eq. (7) and Eq. (8):
${\mathsf{prox}}_{\tau\bar{h}}({\bm{u}})={\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{u}})\quad\forall{\bm{u}}\in\mathbb{R}^{2nd}\;.$ (19)
By Corollary 10.c in [62], there exists a closed and convex $\phi$ such that ${\mathcal{T}}_{2\tau\sqrt{d}}({\bm{u}})\in\partial\phi({\bm{u}})$. Plugging in ${\bm{u}}={\bm{W}}{\bm{z}}$ and multiplying both sides by ${\bm{W}}^{{\mathsf{T}}}$ gives us ${\mathcal{S}}_{\tau}({\bm{z}})\in{\bm{W}}^{{\mathsf{T}}}\partial\phi({\bm{W}}{\bm{z}})$. By the subdifferential chain rule, we know that ${\bm{W}}^{{\mathsf{T}}}\partial\phi({\bm{W}}{\bm{z}})\subset\partial[\phi\circ{\bm{W}}]({\bm{z}})$, where $\circ$ denotes composition of the functions. Therefore, we have
${\mathcal{S}}_{\tau}({\bm{z}})\in\partial[\phi\circ{\bm{W}}]({\bm{z}})\;.$
Next, letting $\psi({\bm{z}})=[\phi\circ{\bm{W}}]({\bm{z}})$, we need to show $\psi$ is closed and convex. Since $\phi$ is closed, $\phi\circ{\bm{W}}$ is also closed. Since $\phi$ is convex and $\bm{W}$} is linear, $\phi\circ{\bm{W}}$ is convex. We can show $\|{\bm{W}}\|_{2}=1$ as follows
$\|{\bm{W}}\|_{2}$ $=\sqrt{\sigma_{\text{max}}({\bm{W}}^{{\mathsf{T}}}{\bm{W}})}$
$=\sqrt{\sigma_{\text{max}}(\bm{I})}=1\;,$ (20)
where $\sigma_{\text{max}}({\bm{I}})$ is the largest eigenvalue of $\bm{I}$}.
We now verify the nonexpansiveness of ${\mathcal{S}}_{\tau}({\bm{z}})$.
$\|{\mathcal{S}}_{\tau}({\bm{z}}_{1})-{\mathcal{S}}_{\tau}({\bm{z}}_{2})\|_{2}$
$\quad=\|{\bm{W}}^{{\mathsf{T}}}{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}}_{1})-{\bm{W}}^{{\mathsf{T}}}{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}}_{2})\|_{2}$
$\quad=\|{\bm{W}}^{{\mathsf{T}}}({\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}}_{1})-{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}}_{2}))\|_{2}$
$\quad\leq\|{\bm{W}}^{{\mathsf{T}}}\|_{2}\|{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}}_{1})-{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}}_{2})\|_{2}$
$\quad=\|{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}}_{1})-{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}}_{2})\|_{2}$
$\quad\leq\|{\bm{W}}{\bm{z}}_{1}-{\bm{W}}{\bm{z}}_{2}\|_{2}$
$\quad\leq\|{\bm{W}}\|_{2}\|{\bm{z}}_{1}-{\bm{z}}_{2}\|_{2}$
$\quad=\|{\bm{z}}_{1}-{\bm{z}}_{2}\|_{2}\;,$
Where the third and last equalities comes from Eq. (20). The second inequality comes from the fact that soft-thresholding is a non-expansive operator. The above inequality is exactly the definition of a nonexpansive operator. ∎
Proof of Proposition 2
Proposition 2.
Let $\widetilde{h}^{a}({\bm{z}})$ and $\widetilde{h}^{i}({\bm{z}})$ denote the Huber variants of anisotropic TV and isotropic TV given in Eq. (9) and Eq. (11), respectively. Then,
${\mathcal{S}}^{a}_{\tau}({\bm{z}})={\bm{z}}-\tau\nabla\widetilde{h}^{a}({\bm{z}})\quad\text{and}\quad{\mathcal{S}}^{i}_{\tau}({\bm{z}})={\bm{z}}-\tau\nabla\widetilde{h}^{i}({\bm{z}}).$
Proof.
We will first consider anisotropic TV ${\mathcal{S}}^{a}_{\tau}({\bm{z}})$. The soft-thresholding operator can be rewritten as
${\mathcal{T}}^{\hskip 1.22911pta}_{\lambda}(u)\,\coloneqq\,{\mathsf{max}}(|u|-\lambda,0)\frac{u}{|u|}=u-\mathop{\mathsf{min}}(\lambda,|u|)\text{sgn}(u).$
Since soft-thresholding only affects the difference components, it can be written as
${\mathcal{S}}_{\tau}({\bm{z}})$ $={\bm{W}}^{{\mathsf{T}}}{\mathcal{T}}^{\hskip 1.22911pta}_{\tau}({\bm{W}}{\bm{z}})$
[…]
$={\bm{z}}-\left[\frac{{\bm{M}}^{{\mathsf{T}}}}{2\sqrt{d}}\frac{{\bm{D}}^{{\mathsf{T}}}}{2\sqrt{d}}\right]\begin{bmatrix}0\\ \vdots\\ 0\\ \mathop{\mathsf{min}}\left(\tau 2\sqrt{d},\frac{|[{\bm{D}}{\bm{z}}]_{1}|}{2\sqrt{d}}\right)\text{sgn}\left(\frac{[{\bm{D}}{\bm{z}}]_{1}}{2\sqrt{d}}\right)\\ \vdots\\ \mathop{\mathsf{min}}\left(\tau 2\sqrt{d},\frac{|[{\bm{D}}{\bm{z}}]_{nd}|}{2\sqrt{d}}\right)\text{sgn}\left(\frac{[{\bm{D}}{\bm{z}}]_{nd}}{2\sqrt{d}}\right)\end{bmatrix}$
$={\bm{z}}-\frac{{\bm{D}}^{{\mathsf{T}}}}{2\sqrt{d}}\begin{bmatrix}\mathop{\mathsf{min}}\left(\tau 2\sqrt{d},\frac{|[{\bm{D}}{\bm{z}}]_{1}|}{2\sqrt{d}}\right)\text{sgn}\left(\frac{[{\bm{D}}{\bm{z}}]_{1}}{2\sqrt{d}}\right)\\ \vdots\\ \mathop{\mathsf{min}}\left(\tau 2\sqrt{d},\frac{|[{\bm{D}}{\bm{z}}]_{nd}|}{2\sqrt{d}}\right)\text{sgn}\left(\frac{[{\bm{D}}{\bm{z}}]_{nd}}{2\sqrt{d}}\right)\end{bmatrix}$
$={\bm{z}}-\frac{{\bm{D}}^{{\mathsf{T}}}}{2\sqrt{d}}\mathop{\mathsf{min}}\left(\tau 2\sqrt{d},\frac{|{\bm{D}}{\bm{z}}|}{2\sqrt{d}}\right)\text{sgn}\left(\frac{{\bm{D}}{\bm{z}}}{2\sqrt{d}}\right)$
$={\bm{z}}-\tau{\bm{D}}^{{\mathsf{T}}}\mathop{\mathsf{min}}\left(1,\frac{|{\bm{D}}{\bm{z}}|}{\tau 4d}\right)\text{sgn}\left({\bm{D}}{\bm{z}}\right).$
In the last two lines, the operations $\mathop{\mathsf{min}}$ and sgn are to be understood as being applied element-wise to ${\bm{D}}{\bm{z}}$. The function $\phi_{\theta}$ from Eq. (10) has the derivative,
$\phi_{\theta}^{\prime}(u)=\begin{cases}\frac{u}{\theta}&\text{for }|u|\leq\theta\\ \text{sgn}(u)&\text{for }|u|>\theta\end{cases}=\mathop{\mathsf{min}}\left(1,\frac{|u|}{\theta}\right)\text{sgn}(u)$
Therefore, the operator can be written as
${\mathcal{S}}^{a}_{\tau}({\bm{z}})={\bm{z}}-\tau{\bm{D}}^{{\mathsf{T}}}\phi^{\prime}_{\tau 4d}\left({\bm{D}}{\bm{z}}\right),$
where $\phi^{\prime}_{\tau 4d}$ is applied element-wise. In other words, ${\mathcal{S}}^{a}_{\tau}({\bm{z}})$ is a gradient step with step-size $\tau$ on
$\widetilde{h}^{a}({\bm{z}})=\sum_{i=1}^{nd}\phi_{\tau 4d}([{\bm{D}}{\bm{z}}]_{i}).$ (21)
A step-size of $\tau$ satisfies the criterion for the step size to be $\frac{1}{L}$, where $L$ is the Lipschitz constant of the gradients. The Lipschitz constant for Eq. 21 is $1/\tau$, which can be verified as follows:
$\|\nabla\widetilde{h}({\bm{x}})-\nabla\widetilde{h}({\bm{y}})\|$ $=\|{\bm{D}}^{{\mathsf{T}}}\phi^{\prime}_{4\tau d}({\bm{D}}{\bm{x}})-{\bm{D}}^{{\mathsf{T}}}\phi^{\prime}_{4\tau d}({\bm{D}}{\bm{y}})\|$
$\leq\|{\bm{D}}\|\|\phi^{\prime}_{4\tau d}({\bm{D}}{\bm{x}})-\phi^{\prime}_{4\tau d}({\bm{D}}{\bm{y}})\|$
$\leq\left(\frac{1}{4\tau d}\right)\|{\bm{D}}\|\|{\bm{D}}{\bm{x}}-{\bm{D}}{\bm{y}}\|$
$\leq\left(\frac{1}{4\tau d}\right)\|{\bm{D}}\|^{2}\|{\bm{x}}-{\bm{y}}\|$
$=\frac{1}{\tau}\|{\bm{x}}-{\bm{y}}\|,$
where we use the fact that the finite difference matrix $\bm{D}$} has spectral norm $2\sqrt{d}$.
Now consider isotropic ${\mathcal{S}}^{i}_{\tau}({\bm{z}})$. The soft-thresholding operator can be rewritten as
${\mathcal{T}}^{\hskip 1.22911pt\text{i}}_{\lambda}({\bm{u}}):={\mathsf{max}}(\|{\bm{u}}\|_{2}-\lambda,0)\frac{{\bm{u}}}{\|{\bm{u}}\|_{2}}={\bm{u}}-\mathop{\mathsf{min}}(\lambda,\|{\bm{u}}\|)\frac{{\bm{u}}}{\|{\bm{u}}\|}.$
Let $G_{i}$ for $i\in\{1,...,n\}$ be the set of $d$ indices corresponding to the differences along all $d$ dimensions for $x_{i}$. Since soft-thresholding affects only the difference components, it can be rewritten
${\mathcal{S}}_{\tau}({\bm{z}})$ $={\bm{W}}^{{\mathsf{T}}}{\mathcal{T}}^{\hskip 1.22911pti}_{\tau}({\bm{W}}{\bm{z}})$
$={\bm{z}}-\frac{{\bm{D}}^{{\mathsf{T}}}}{2\sqrt{d}}\begin{bmatrix}\mathop{\mathsf{min}}\left(\tau 2\sqrt{d},\frac{\|[{\bm{D}}{\bm{z}}]_{G_{1}}\|}{2\sqrt{d}}\right)\frac{[{\bm{D}}{\bm{z}}]_{G_{1}}}{\|[{\bm{D}}{\bm{z}}]_{G_{1}}\|}\\ \vdots\\ \mathop{\mathsf{min}}\left(\tau 2\sqrt{d},\frac{\|[{\bm{D}}{\bm{z}}]_{G_{n}}\|}{2\sqrt{d}}\right)\frac{[{\bm{D}}{\bm{z}}]_{G_{n}}}{\|[{\bm{D}}{\bm{z}}]_{G_{n}}\|}\end{bmatrix}$
$={\bm{z}}-\tau{\bm{D}}^{{\mathsf{T}}}\mathop{\mathsf{min}}\left(1,\frac{\|[{\bm{D}}{\bm{z}}]_{G_{i}}\|}{\tau 4d}\right)\frac{[{\bm{D}}{\bm{z}}]_{G_{i}}}{\|[{\bm{D}}{\bm{z}}]_{G_{i}}\|}.$
The last line is meant to be understood as applying $\mathop{\mathsf{min}}$ and the norm to each grouping of indices. The function $\psi_{\theta}$ from Eq. (12) has the gradient
$\nabla\psi_{\theta}({\bm{u}})=\begin{cases}\frac{{\bm{u}}}{\theta}&\text{for }\|{\bm{u}}\|\leq\theta\\ \frac{{\bm{u}}}{\|{\bm{u}}\|}&\text{for }\|{\bm{u}}\|>\theta\end{cases}=\mathop{\mathsf{min}}\left(1,\frac{\|{\bm{u}}\|}{\theta}\right)\frac{{\bm{u}}}{\|{\bm{u}}\|}$
and so the operator can be written as
${\mathcal{S}}^{i}_{\tau}({\bm{z}})={\bm{z}}-\tau{\bm{D}}^{{\mathsf{T}}}\nabla\phi_{\tau 4d}({\bm{D}}{\bm{z}}),$
where $\nabla\phi_{\tau 4d}$ is applied to each grouping of indices. That is, ${\mathcal{S}}^{i}_{\tau}({\bm{z}})$ is a gradient step with the step-size $\tau$ on
$\widetilde{h}^{i}({\bm{z}})=\sum_{i=1}^{n}\phi_{\tau 4d}([{\bm{D}}{\bm{z}}]_{G_{i}}).$ (22)
Again, $\tau$ satisfies the step-size to be $\frac{1}{L}$, since the Lipschitz constant of the gradients for Eq. 22 remains $1/\tau$. ∎
Proof of Proposition 3
Proposition 3.
For all ${\bm{z}}\in\mathbb{R}^{n}$, the operator ${\mathcal{S}}_{\tau}({\bm{z}})$ satisfies
  1. (a)
    ${\mathcal{S}}_{\tau}({\bm{z}})={\mathsf{prox}}_{\tau{h}}({\bm{z}}+{\bm{\delta}})$ where $\|{\bm{\delta}}\|\leq\tau\epsilon_{1}$,
  2. (b)
    ${\bm{z}}-{\mathcal{S}}_{\tau}({\bm{z}})\in\tau\partial_{\tau\epsilon_{2}}{h}({\mathcal{S}}_{\tau}({\bm{z}}))$,
where $\epsilon_{1}$ and $\epsilon_{2}$ are constants.
Proof.
Proof of Part (a). We know $\|{\bm{W}}\|_{2}=1$ from (20). For both the anisotropic and isotropic versions, by its definition the soft-thresholding operator ${\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})$ at most changes each of the $nd$ difference elements of ${\bm{W}}{\bm{z}}$ by $2\tau\sqrt{d}$, so we can bound
$\|{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})-{\bm{W}}{\bm{z}}\|_{2}$ $\leq\sqrt{nd(2\tau\sqrt{d})^{2}}$
$=\sqrt{4\tau^{2}nd^{2}}$
$=2\tau d\sqrt{n}\;.$ (23)
We bound $\|{\mathcal{S}}_{\tau}({\bm{z}})-{\bm{z}}\|_{2}$ by Eq. (20) and Eq. (23)
$\Big{\|}{\mathcal{S}}_{\tau}({\bm{z}})-{\bm{z}}\Big{\|}_{2}$ $=\Big{\|}{\mathcal{S}}_{\tau}({\bm{z}})-{\bm{W}}^{{\mathsf{T}}}{\bm{W}}{\bm{z}}\Big{\|}_{2}$
$\leq\Big{\|}{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})-{\bm{W}}{\bm{z}}\Big{\|}_{2}$
$\leq 2\tau d\sqrt{n}\;.$ (24)
For $\Phi_{\tau}({\bm{y}}):=\frac{1}{2\tau}\|{\bm{y}}-{\bm{z}}\|_{2}^{2}+{h}({\bm{y}})$, its subdifferential at ${\mathcal{S}}_{\tau}({\bm{z}})$ is
$\partial\Phi_{\tau}({\mathcal{S}}_{\tau}({\bm{z}}))$ $=\frac{1}{\tau}({\mathcal{S}}_{\tau}({\bm{z}})-{\bm{z}})+\partial{h}({\mathcal{S}}_{\tau}({\bm{z}}))$
$=\frac{1}{\tau}({\mathcal{S}}_{\tau}({\bm{z}})-{\bm{z}})+\partial[\bar{h}\circ{\bm{W}}]({\mathcal{S}}_{\tau}({\bm{z}}))$
$=\frac{1}{\tau}({\mathcal{S}}_{\tau}({\bm{z}})-{\bm{z}})+{\bm{W}}^{{\mathsf{T}}}\partial\bar{h}({\bm{W}}{\mathcal{S}}_{\tau}({\bm{z}}))\;,$
where the second equality comes from $\bar{h}\circ{\bm{W}}={h}$ and the third equality comes from the chain rule.
The distance $d(\bm{0},\partial\Phi_{\tau}({\mathcal{S}}_{\tau}({\bm{z}}))$ is defined as the minimum $\ell_{2}$ distance between $\bm{0}$ and any vector in $\partial\Phi_{\tau}({\mathcal{S}}_{\tau}({\bm{z}}))$. Therefore, letting ${\bm{g}}_{1}\in\partial\bar{h}({\bm{W}}{\mathcal{S}}_{\tau}({\bm{z}}))$, we get the following inequality
$d(\bm{0},\partial\Phi_{\tau}({\mathcal{S}}_{\tau}({\bm{z}}))$ $\leq\Big{\|}\bm{0}-\left(\frac{1}{\tau}({\mathcal{S}}_{\tau}({\bm{z}})-{\bm{z}})+{\bm{W}}^{{\mathsf{T}}}{\bm{g}}_{1}\right)\Big{\|}_{2}$
$\leq\frac{1}{\tau}\Big{\|}{\mathcal{S}}_{\tau}({\bm{z}})-{\bm{z}}\Big{\|}_{2}+\Big{\|}{\bm{W}}^{{\mathsf{T}}}{\bm{g}}_{1}\Big{\|}_{2}$
$\leq\frac{1}{\tau}(2\tau d\sqrt{n})+\|{\bm{g}}_{1}\|_{2}$
$\leq\frac{1}{\tau}(2\tau d\sqrt{n})+2d\sqrt{n}=\frac{\tau\epsilon_{1}}{\tau}\;,$ (25)
with $\epsilon_{1}=4d\sqrt{n}$. The third inequality comes from (24) and nonexpansiveness of ${\bm{W}}^{{\mathsf{T}}}$. The fourth inequality comes from the fact that the norm of any gradient of $\bar{h}$ is bounded by $2d\sqrt{n}$, as shown in Lemma 1.
Letting ${\bm{g}}_{2}\in\partial\Phi_{\tau}({\mathcal{S}}_{\tau}({\bm{z}}))$, we can write $\|{\bm{g}}_{2}\|\leq\frac{\tau\epsilon_{1}}{\tau}$. Equivalently,
${\bm{g}}_{2}\in\frac{1}{\tau}({\mathcal{S}}_{\tau}({\bm{z}})-{\bm{z}})+\partial{h}({\mathcal{S}}_{\tau}({\bm{z}}))$
$\iff$
${\bm{z}}+{\bm{\delta}}-{\mathcal{S}}_{\tau}({\bm{z}})\in\tau\partial{h}({\mathcal{S}}_{\tau}({\bm{z}}))$
$\iff$
${\mathcal{S}}_{\tau}({\bm{z}})={\mathsf{prox}}_{\tau{h}}({\bm{z}}+{\bm{\delta}}),\text{ where }\|{\bm{\delta}}\|_{2}\leq\tau\epsilon_{1}\;,$
and ${\bm{\delta}}=\tau{\bm{g}}_{2}$. The last equivalence comes from the property of proximal operators that ${\bm{x}}={\mathsf{prox}}_{\tau{h}}({\bm{y}})\iff{\bm{y}}-{\bm{x}}\in\tau\partial\hskip 1.49994pt{h}({\bm{x}})$. This establishes the desired result of part (b).
Proof of Part (b). From the definition of the subgradient and the proximal operator, we have
${\bm{u}}={\mathsf{prox}}_{\tau\bar{h}}({\bm{W}}{\bm{z}})\quad\iff\quad{\bm{W}}{\bm{z}}-{\bm{u}}\in\tau\partial\bar{h}({\bm{u}})\;.$
By definition of subgradient, $\forall{\bm{w}}\in\mathbb{R}^{2nd}$
$\bar{h}({\bm{w}})$ $\geq\bar{h}({\bm{u}})+\left(\frac{{\bm{W}}{\bm{z}}-{\bm{u}}}{\tau}\right)^{{\mathsf{T}}}({\bm{w}}-{\bm{u}})\;.$
By replacing $\bm{w}$} with ${\bm{W}}{\bm{y}}$, where ${\bm{y}}\in\mathbb{R}^{n}$, we obtain
$\bar{h}({\bm{W}}{\bm{y}})$ $\geq\bar{h}({\bm{u}})+\frac{1}{\tau}({\bm{W}}{\bm{z}}-{\bm{u}})^{{\mathsf{T}}}({\bm{W}}{\bm{y}}-{\bm{u}})$
$=\bar{h}({\bm{u}})+\frac{1}{\tau}\Big{(}({\bm{W}}{\bm{z}})^{{\mathsf{T}}}({\bm{W}}{\bm{y}})-({\bm{W}}({\bm{z}}+{\bm{y}}))^{{\mathsf{T}}}{\bm{u}}+\|{\bm{u}}\|_{2}^{2}\Big{)}$
$\geq\bar{h}({\bm{u}})+\frac{1}{\tau}\Big{(}{\bm{z}}^{{\mathsf{T}}}{\bm{y}}-({\bm{z}}+{\bm{y}})^{{\mathsf{T}}}({\bm{W}}^{{\mathsf{T}}}{\bm{u}})+\|{\bm{W}}^{{\mathsf{T}}}{\bm{u}}\|_{2}^{2}\Big{)}$
$=\bar{h}({\bm{u}})+\frac{1}{\tau}({\bm{z}}-{\bm{W}}^{{\mathsf{T}}}{\bm{u}})^{{\mathsf{T}}}({\bm{y}}-{\bm{W}}^{{\mathsf{T}}}{\bm{u}})$
$=\bar{h}({\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}}))+\frac{1}{\tau}({\bm{z}}-{\mathcal{S}}_{\tau}({\bm{z}}))^{{\mathsf{T}}}({\bm{y}}-{\mathcal{S}}_{\tau}({\bm{z}}))\;,$ (26)
where the last inequality comes from the definition of adjoint $({\bm{W}}{\bm{x}}_{1})^{{\mathsf{T}}}{\bm{x}}_{2}={\bm{x}}_{1}^{{\mathsf{T}}}({\bm{W}}^{{\mathsf{T}}}{\bm{x}}_{2})$, the fact that ${\bm{W}}^{{\mathsf{T}}}{\bm{W}}={\bm{I}}$, and that $\|{\bm{W}}^{T}\|_{2}=1$. In the last line, we used the fact that ${\bm{u}}={\mathsf{prox}}_{\tau\bar{h}}({\bm{W}}{\bm{z}})={\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})$ from Eq. 19 and ${\mathcal{S}}_{\tau}({\bm{z}})={\bm{W}}^{{\mathsf{T}}}{\bm{u}}$ .
Figure 5: A visual illustration of the characterization of the projection as ${\bm{W}}{\bm{W}}^{{\mathsf{T}}}{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})={\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})+{\bm{\delta}}$. For any ${\bm{z}}\in\mathbb{R}^{n}$, the vector ${\bm{W}}{\bm{z}}\in{\mathcal{U}}:=\{{\bm{u}}|\ {\bm{u}}={\bm{W}}{\bm{W}}^{{\mathsf{T}}}{\bm{u}}\}$. Applying the soft-thresholding function gives ${\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})$, not necessarily in $\mathcal{U}$}. Applying ${\bm{W}}{\bm{W}}^{{\mathsf{T}}}$ projects back to $\mathcal{U}$}, which can be represented as ${\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})+{\bm{\delta}}$.
The projection can be characterized as
${\bm{W}}{\bm{W}}^{{\mathsf{T}}}{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})={\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})+{\bm{\beta}}\;.$ (27)
That is, $\bm{\beta}$} is the orthogonal component of ${\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})$ to the subspace ${\mathcal{U}}:=\{{\bm{u}}|\ {\bm{u}}={\bm{W}}{\bm{W}}^{{\mathsf{T}}}{\bm{u}}\}$. Figure 5 provides a visual representation.
Since the projection ${\bm{W}}{\bm{W}}^{{\mathsf{T}}}$ minimizes the $\ell_{2}$ distance between ${\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})$ and $\mathcal{U}$}, the following inequality holds
$\|{\bm{\beta}}\|_{2}$ $=\mathop{\mathsf{min}}_{{\bm{w}}\in{\mathcal{U}}}\|{\bm{w}}-{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})\|_{2}$
$\leq\|{\bm{w}}-{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})\|_{2}\qquad\forall{\bm{w}}\in{\mathcal{U}}\;.$
Since ${\bm{W}}{\bm{z}}\in{\mathcal{U}}$, this yields
$\|{\bm{\beta}}\|_{2}\leq\|{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})-{\bm{W}}{\bm{z}}\|_{2}\leq 2\tau d\sqrt{n}\;,$
where the last inequality comes from (23). Letting ${\bm{\beta}}^{\text{dif}}$ be the components of $\bm{\delta}$} corresponding to difference coefficients,
$\|{\bm{\beta}}^{\text{dif}}\|_{2,1}\leq\|{\bm{\beta}}^{\text{dif}}\|_{1,1}$ $=\|{\bm{\beta}}^{\text{dif}}\|_{1}$
$\leq\sqrt{nd}\|{\bm{\beta}}^{\text{dif}}\|_{2}$
$\leq\sqrt{nd}\|{\bm{\beta}}\|_{2}$
$\leq(\sqrt{nd})(2\tau d\sqrt{n})$
$=2\tau nd^{3/2}\;,$
where the first inequality comes from the fact that the $\ell_{2}$ norm is smaller than the $\ell_{1}$ norm for any given vector. The second inequality comes from the fact that the $\ell_{1}$ norm is at most the square root of the dimension of the vector times its $\ell_{2}$ norm.
Using the definition of $\bar{h}$ in Eq. (16) relationship $\bar{h}({\bm{W}}{\bm{z}})={h}({\bm{z}})$, shown in Eq. (18), we get
${h}({\mathcal{S}}_{\tau}({\bm{z}}))$ $=\bar{h}({\bm{W}}{\mathcal{S}}_{\tau}({\bm{z}}))$
$=2\sqrt{d}\|[{\bm{W}}{\mathcal{S}}_{\tau}({\bm{z}})]^{\text{dif}}\|_{p,1}$
$=2\sqrt{d}\|[{\bm{W}}{\bm{W}}^{{\mathsf{T}}}{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})]^{\text{dif}}\|_{p,1}$
$=2\sqrt{d}\|[{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})+{\bm{\beta}}]^{\text{dif}}\|_{p,1}$
$\leq 2\sqrt{d}\|[{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})]^{\text{dif}}\|_{p,1}$
$\qquad\qquad\qquad+2\sqrt{d}\|{\bm{\beta}}^{\text{dif}}\|_{p,1}$
$\leq 2\sqrt{d}\|[{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}})]^{\text{dif}}\|_{p,1}$
$\qquad\qquad\qquad+(2\sqrt{d})(2\tau nd^{3/2})$
$=\bar{h}({\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}}))+4\tau nd^{2}\;.$
Rearranging the above inequality yields
$\bar{h}({\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}}))\geq{h}({\mathcal{S}}_{\tau}({\bm{z}}))-4\tau nd^{2}\;.$ (28)
From Eq. (26) we have
${h}({\bm{y}})\geq\bar{h}({\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}}))+\frac{1}{\tau}({\bm{z}}-{\mathcal{S}}_{\tau}({\bm{z}}))^{{\mathsf{T}}}({\bm{y}}-{\mathcal{S}}_{\tau}({\bm{z}}))\;.$
Combining this with Eq. (28) provides the following for all ${\bm{y}}\in\mathbb{R}^{n}$
${h}({\bm{y}})\geq{h}({\mathcal{S}}_{\tau}({\bm{z}}))+\frac{1}{\tau}({\bm{z}}-{\mathcal{S}}_{\tau}({\bm{z}}))^{{\mathsf{T}}}({\bm{y}}-{\mathcal{S}}_{\tau}({\bm{z}}))-\tau 4nd^{2}\;.$
By Definition 2 this establishes
${\bm{z}}-{\mathcal{S}}_{\tau}({\bm{z}})\in\tau\partial_{\tau\epsilon_{2}}{h}({\mathcal{S}}_{\tau}({\bm{z}}))\;,$
where $\epsilon_{2}=4nd^{2}$. This establishes the desired result for Part (c). ∎
Lemma 1.
For both the anisotropic and isotropic versions, any subgradient ${\bm{g}}\in\partial\bar{h}({\bm{u}})$ has a bounded $\ell_{2}$ norm $\|{\bm{g}}\|_{2}\leq 2d\sqrt{n}$.
Proof.
This lemma makes use of the following subdifferential sum property: for any functions $f_{1},\dots,f_{m}\in\Gamma^{0}$ such that the sets $\text{ri}(\text{dom}f_{i})$ have a point in common, where ri is the relative interior, then
$\partial\Bigg{[}\sum_{i=1}^{m}f_{i}\Bigg{]}({\bm{x}})=\sum_{i=1}^{m}\partial f_{i}({\bm{x}})\qquad\forall{\bm{x}}\in\mathbb{R}^{n}\;.$ (29)
See Theorem 23.8 in [66] and Corollary 16.48 in [70]. First consider the anisotropic version
$\bar{h}^{\text{a}}({\bm{u}})=2\sqrt{d}\|{\bm{u}}^{\text{dif}}\|_{1,1}=2\sqrt{d}\sum_{i=1}^{n}\|[{\bm{u}}^{\text{dif}}]_{i}\|_{1}\;.$
Since the domain of each $||\cdot||_{1}$ is all of $\mathbb{R}^{2nd}$, the relative interiors of the domains have a nonempty intersection, and so we can apply Eq. (29). Consider $\partial\big{(}\|[{\bm{u}}^{\text{dif}}]_{i}\|_{1}\big{)}$, whose norm is bounded by $\sqrt{d}$. By Eq. (29), letting ${\bm{g}}\in\partial\bar{h}^{\text{a}}({\bm{u}})$, we can bound
$\|{\bm{g}}\|_{2}$ $=2\sqrt{d}\sqrt{\sum_{i=1}^{n}\Big{(}\partial\big{(}\|[{\bm{u}}^{\text{dif}}]_{i}\|_{1}\big{)}\Big{)}^{2}}$
$\leq(2\sqrt{d})\sqrt{n(\sqrt{d})^{2}}$
$=2d\sqrt{n}\;.$
Now consider the isotropic version
$\bar{h}^{\text{i}}({\bm{u}})=2\sqrt{d}\|{\bm{u}}^{\text{dif}}\|_{2,1}=2\sqrt{d}\sum_{i=1}^{n}\|[{\bm{u}}^{\text{dif}}]_{i}\|_{2}\;.$
Similarly, the domain of each $||\cdot||_{2}$ is all of $\mathbb{R}^{2nd}$, and so we can apply Eq. (29). Consider $\partial\big{(}\|[{\bm{u}}^{\text{dif}}]_{n}\|_{2}\big{)}$, whose norm is bounded by $1$. By Eq. (29), letting ${\bm{g}}\in\partial\bar{h}^{\text{i}}({\bm{u}})$, we obtain
$\|{\bm{g}}\|_{2}=2\sqrt{d}\sqrt{\sum_{i=1}^{n}\Big{\|}\partial\Big{(}\|[{\bm{u}}^{\text{dif}}]_{i}\|_{2}\Big{)}\Big{\|}_{2}^{2}}\leq 2\sqrt{nd}\;.$
Since $d\geq 1$, we have $2\sqrt{nd}\leq 2d\sqrt{n}$. Therefore, for both the anisotropic and isotropic, $\|{\bm{g}}\|_{2}$ is bounded by $2d\sqrt{n}$. ∎